A small hospital has 130 full-time employees. Twenty percent are out with the Flu. How many employees are sick and how many are healthy? sick#_____ healthy#_______
Answer #1

ANSWER.

* Calculation of number of employees sick and healthy.

            Total number of emplyees = 130 .

            Employees exposed with the flu = 20 % (20/100)

            Employees not exposed with flu = 100% - 20%

                                                                = 80% (80/100).

    So

        Number of sick employees = 130 * 20 / 100

                                                    = 2600 / 100

                                                    = 26.

       Number of healthy employees = 130 * 80 / 100

                                                        = 10400 / 100

                                                        = 104.